OpenStudy (anonymous):

If g(x) = .5x + 3, find the value of g(0) + 2. A. 5 B. 4 C. 5.5 D. 7

OpenStudy (anonymous):

so g(0) = .5x0 + 3 = 3 . g(0) + 2 = 3+2 = 5 i.e. A